# Good News for Polynomial Root-finding

### v35

@@ 19.5 Extension to counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 19.3 that up to a constant factor the overall number of these suspect squares is given by their number at the partition step. Such a result will extend to all subdivision steps our count $n^{\prime}=O(m)$ of (19.3), so far restricted to partition steps. Together with Lemma 19.1 this will complete our proof of Thm. 19.1.
@@
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 17.2.)
@@
The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic. E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 19.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares. (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares. (iii) In the proof of Lemma 19.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$ but the bounds become stronger as $n_{i}$ increases. (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 19.1 can be strengthened at least in this case.
@@
Alg. 18.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm. 19.1 only involves $\eta$ the argument in Sec. 18.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
20 Computation of a point near root set and root-counting in a disc
@@ 20.1 Computation of a point near root set: introductory comments
Recall Schröder’s iterations [128], celebrated for fast convergence to a root of multiplicity $m$ [77, Sec. 5.4]:
$z=y-m~{}{\rm NR}(y).$ (20.1)
Due to map (1.4) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ zeros of $p$ and readily verify
Observation 20.1.
Schröder’s single iteration (20.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the zero $z_{1}$ of $p(x)$ unless $y=z_{1}$.
Next we extend this observation by proving that the output $z$ of Schröder’s iteration (20.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contain all $d$ zeros of $p$, and point $y$ lies far enough from the disc $D$. Here are our specific estimates.
@@ 20.2 The case where a disc contains all $d$ roots
Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (20.1) for $m=d$. Let $\bar{x}\in CH$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}$.
@@
Under the assumptions of Thm. 20.2 it holds that $r_{d}(z)\leq 2.25\Delta$ for $|y|\geq 4$, while $r_{d}(z)\leq 25\Delta/16$ for $|y|\geq 6$.
Readily verify
Observation 20.2.
Let #$\#(D)=m$ for a disc $D$ and let $\Delta:=\Delta(CH(X(D)))$ denote the diameter of the convex hull of the root set $X(D)$ in that disc. Write $r_{d-m}(z,p):=0$, $r:=r_{d-m+1}(z,p)$, and $\rho:=r_{d}(z,p)$ for a complex $z$ and let $r+\Delta\leq r_{d-m}(z,p)$, which implies that $r$ and $\rho$ do not depend on the $d-m$ zeros of $p$ lying outside the disc $D$. Then $r-\rho\leq\Delta\leq 2r$.
Corollary 20.2.
If $\rho\leq\nu\Delta$ under the assumptions of Observation 20.2, then the disc $D(z,r)$ is $1/(2+2\nu)$-rigid.
Now we devise our fast compression algorithm.
@@
OUTPUT: a complex $z$, and a positive $r$ such that (i) $X(D(z,r))=X(D(0,1))$ r equivalently #$\#(D(z,r))=d$, and (ii) $r\leq\epsilon$ or the disc $D(z,r)$ is rigid.
@@
(i) Compute and output $z=y-d{\rm NR}~{}(y)$ (cf. (20.1)).
(ii) Compute an approximation $\tilde{r}\geq 0$ by applying Alg. 5.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$).
@@
Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into the rigid disc $D(z,r)$.
Correctness of the algorithm is readily verified if Alg. 5.1 outputs a range $[\rho_{-}\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta>1$ of Alg. 5.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\nu+2}$-rigid by virtue of Cor. 20.2 where $\nu\leq 2.25$ by virtue of Cor. 20.1 since $y>4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is rigid.
This compression queries the evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 5.1, estimated in Cor. 5.1. We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$. Then we yield desired compression if the algorithm outputs $r$ exceeding 1/4 of that tolerance value. Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
@@ 20.3 Extension to any well-isolated disc
Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$. Then verify that the $d-m$ external zeros of $p$, lying outside the disc $D(0,1)$, little affect the proof of Thm. 20.1 and our whole study in the previous subsection.
Thus we first perform $O(\log(d))$ subdivision steps (see Observation 17.3). We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc.
Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
20.4 A fast root-counter based on Schröder’s iterations
@@
Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we arrive at a root-counter in $D(0,1)$ that evaluates $p(x)$ at two points.
@@
In this case $|m{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
If #$\#(D(0,1))=\bar{m}\neq m$, however, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be positive if $|y|$ is sufficiently (although reasonably) large.
@@ 20.5 Overall complexity of our root-finders
We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests. We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (each included to yield isolation of an input disc for compression and reduced to $O(m)$ e/i tests), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(d))$ e/i tests for approximation of the root radius $r_{d-m+1}(z,p)$.
Summarize these estimates, combine them with the estimates of the previous subsection, and arrive at the complexity bounds of our Main Theorems.
Remark 20.1.
If we incorporate Kirrinnis’s algorithm for multipoint polynomial evaluation, [65, Alg. 5.3], rather than Moroz’s algorithm of [78], then our root-finders would slow down but still would stay very fast and still would use near optimal number of bit operations in the cases where $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model. Indeed, by using our compression steps we decrease the complexity estimate of (8.4) and (15.4) by a factor of $b/\log(b)$ to arrive at the bound $\mathbb{B}_{\rm roots}^{\prime}=\tilde{O}(((m^{2}+d)(m^{2}+d+b))$ on the number of bit operations. This bound turns into $\tilde{O}((d+b)d)$ if $m^{2}=\tilde{O}(d)$ and whp under the Random Root Model.
APPENDIX
Appendix A Applications and implementation of root-finders: brief overview
Numerous areas of mathematics and computational mathematics studied in the last 100 years had no direct link to polynomial root-finding. Furthermore, many computational problems arising in the sciences, engineering, business management, and statistics have been linearized and then solved by using tools from linear algebra, linear programming, fast Fourier transform (FFT), and Fast Multipole Method (FMM). [^44] Both FFT and FMM are among the Top Ten Algorithms of 20-th century (cf. [26]). For the history of FFT see D.E. Knuth [66]; on FMM see L.F. Greengard and V. Rokhlin Jr. [48] and [147]. Such tools sometimes involve root-finding for polynomials but usually of smaller degree $d$, where the available subroutines are sufficiently effective in most cases.
Root-finding for large degree polynomials, however, still has important applications to various areas of Scientific Computing, in particular to Computer Algebra, signal processing, geometric modeling, control, financial mathematics, and complex dynamics (see McNamee and Pan [79, Introduction]).
Unlike the near optimal polynomial root-finders of [92, 99], various non-optimal complex polynomial root-finders have been implemented (see e.g., MATLAB “roots”, Maple “solve”, MPSolve, and CCluster). In particular, Imbach, P., and Yap have implemented the root-finder of [22] in [61].
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [41], but the later version of MPSolve [21] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Randig, Schleicher, and R. Stoll in [126] and by Schleicher and R. Stoll in [138] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see our Sec. D.4).
The root-finder of [22], efficiently implemented in [61], slightly outperforms MPSolve for root-finding in a region of the complex plane containing a small number of roots, although it is by far inferior for the approximation of all $d$ roots of a polynomial $p(x)$. [^45] The computational cost of root-finders of [92, 99, 22, 23] decreases at least proportionally to the number of roots in a region of interest such as a disc or a rectangle, while Ehrlich’s, Weierstrass’s, and a number of other functional iterations approximate the roots in such a region almost as slow as all complex roots. [57, 60] have already superseded MPSolve according to extensive tests for standard test polynomials, and our present root-finders are significantly and probably even dramatically faster (see Sec. 1.9).
Real polynomial root-finding is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest and typically are much less numerous than all $d$ complex roots (cf. Kac [63], Edelman and Kostlan [37]). Current user’s choice for this task is the root-finder in [67], although it has been already superseded in [59] according to extensive tests for standard test polynomials. Our novel algorithm of Sec. 9 can be readily implemented and seems to be the first that efficiently solves the important numerical problem of approximation of both real and near real zeros of a polynomial.
@@ Appendix A Applications and implementation of root-finders: brief overview
Appendix B Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
B.1 Polynomial Factorization and its links to Root-finding (the results)
@@
The root-finders of [130, 92, 99] solve Problems 0 and $*{}^{*}$ by extending the the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^46] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark C.4 and Sec. 7) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [145, 17, 8, 33, 34, 142]. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (B.1)
@@
Definition B.1.
Polynomials with bounded cluster sizes. The zero set of $p$ consists of clusters, made up of a constant number of zeros, each having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
@@
In Sec. B.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [130, Cor. 4.3 and Sec. 19]:
Theorem B.2.
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. B.1, with bounded cluster sizes.
By combining Thms. B.1 and B.2 obtain the following estimates.
Theorem B.3.
@@
(ii) apply the algorithm that supports Thm. B.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (B.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.474747We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
B.2 Lower bounds on the bit operation complexity

### v36

@@ 22.5 Extension to counting all e/i steps
Next consider a non-compact component, count all suspect squares processed in all subdivision steps applied until it is partitioned, and then prove in Lemma 22.3 that up to a constant factor the overall number of these suspect squares is bounded by their overall number at all partition steps. Such a result extends to all subdivision steps the bound $n^{\prime}=O(m)$ of (22.3), so far restricted to partition steps. Together with Lemma 22.1 this will complete our proof of Thm. 22.1.
@@
Let $k+1$ successive subdivision steps with $\sigma$-soft e/i tests be applied to a non-compact component for $\sigma$ of our choice in the range $1<\sigma<\sqrt{2}$. Let the $j$-th step output $n_{j}$ non-adjacent suspect squares for $j=0,1,\dots,k$. Then $\sum_{j=0}^{k-1}n_{j}\leq n_{k}\frac{10\sqrt{2}}{a}$ for the separation fraction $a:=2-\sigma\sqrt{2}$ (cf. Observation 20.1.)
@@
The overhead constants hidden in the "O" notation of our estimates are fairly large but overly pessimistic. E.g., (i) the roots $x$ and $y$ involved in the proof of Lemma 22.2 are linked by a path of suspect squares, but we deduce the estimates of the lemma by using only a part of this path, namely, the part that covers the gaps between non-adjacent suspect squares. (ii) Our proof of that lemma covers the worst case where a long chains between the roots $x$ and $y$ can consist of up to $4m$ suspect squares, but for typical inputs such chains are much shorter and can consist of just a few suspect squares. (iii) In the proof of Lemma 22.3 we assume that $n_{i}-1\geq n_{i}/2$; this holds for any $n_{i}\geq 2$, but the bounds are strengthened as $n_{i}$ increases. (iv) In a component made up of five suspect squares at least two are non-adjacent, and so the bound of Lemma 22.1 can be strengthened at least in this case.
@@
Alg. 21.1 compresses an input disc $D(0,1)$ into $\eta$-rigid disc for $\eta=5/8$, but our proof of Thm. 22.1 only involves $\eta$ in the argument in Sec. 21.4, and this argument and therefore the whole proof of the theorem can be immediately extended if it relies on compression into any rigid disc.
23 A point near root set and root-count in a disc
We prove that the algorithms of this section work where an input disc is $d^{\nu}$-isolated for a fixed positive constant $\nu$. In application to compression we begin with an isolated disc, say, 3.5-isolated, but can turn it into $d^{\nu}$-isolated for any fixed positive constant $\nu$ by applying $h$ additional subdivision steps or $h$-lifting for $h=O(\log(d))$. This comment can be extended to application of e/i test of Sec. 23.4 instead of Cauchy-based e/i tests of Part II.
@@ 23.1 Computation of a point near root set: introductory comments
Recall Schröder’s iterations [132],
$z=y-m~{}{\rm NR}(y),$ (23.1)
which fast converge to a root of multiplicity $m$ [80, Sec. 5.4]. Due to map (1.3) we only need to consider the case where the input disc $D$ is the unit disc $D(0,1)$. We first assume that it contains all $d$ roots and readily verify
Observation 23.1.
Schröder’s single iteration (23.1) for $m=d$ and $p(x)=p_{d}(x-z_{1})^{d}$ outputs the root $z_{1}$ unless $y=z_{1}$.
Next we extend this observation by proving that the output $z$ of Schröder’s iteration (23.1) lies near the convex hull $CH(X(D))$ provided that $m=d$, a disc $D$ contains all $d$ roots, and point $y$ lies far enough from the disc $D$. Here are our specific estimates.
@@ 23.2 The case where a disc contains all $d$ roots
Write $D:=D(0,1)$, $CH:=CH(D)$, and $\Delta:=\Delta(D)$, fix $y$ such that $|y|>2$, let $z:=y-d~{}{\rm NR(y)}$, which is (23.1) for $m=d$. Let $\bar{x}\in CH$. Then $|z-\bar{x}|\leq\alpha~{}\Delta$ for $\alpha:=\frac{(|y|-1)^{2}}{(y|-2)^{2}}=(1+\frac{1}{(y|-2})^{2}$, and so $\alpha=2.25$ for $u|=4$ and $\alpha=25/16$ for $u|=6$, $\alpha\mapsto 1$ $|y|\mapsto\infty$.
@@
Under the assumptions of Thm. 23.2 the disc $D(z,r_{1}(z))$ is $\frac{1}{2+2\alpha}$-rigid.
Proof.
Verify the claim in the case where $x_{d}=0$, $x_{1}=\Delta=1$, e.g., $0\leq x_{j}\leq 1$ for all $j$, and $z=1+\alpha$ and notice that this is a worst case input. ∎
Now we devise our fast compression algorithm inthe case where $m=d$.
@@
OUTPUT: a complex $z$, and a positive $r$ such that #$\#(D(z,r))=d$ and either $r\leq\epsilon$ or otherwise the disc $D(z,r)$ is $\frac{2}{13}$-rigid.
@@
(i) Compute and output $z=y-d~{}{\rm NR}~{}(y)$.
(ii) Compute an approximation $\tilde{r}\geq 0$ to $r_{1}(z)$ by applying Alg. 9.1 for $m=d$, a constant $\beta>1$, and $\epsilon$ replaced by $\epsilon/4$.
@@
Otherwise output $r=\tilde{r}+0.25\epsilon$, defining compression into rigid disc $D(z,r)$.
Correctness of the algorithm is readily verified if Alg. 9.1 outputs a range $[\rho_{-},\rho_{+}]$ of length at most $\epsilon/4$, for in this case $r_{1}\leq r=\tilde{r}+\epsilon/4\leq\epsilon$, and #$\#(D(z,\epsilon))=d$.
Otherwise $1/\beta\leq\tilde{r}/r_{1}\leq\beta$ for a constant $\beta\geq 1$ of Alg. 9.1, while the disc $D(x,r_{1})$ is $\frac{1}{2\alpha+2}$-rigid by virtue of Cor. 23.1 where $\alpha\leq 2.25$ by virtue of THm. 23.1 because $y\geq 4$ in our case. It follows that the disc $\bar{D}:=D(z,r_{1}(z))$ is $\frac{2}{13}$-rigid.
Remark 23.1.
By the same argument we can ensure that unless $\tilde{r}\leq 3\epsilon/4$ the disc $D(z,r_{1}(z))$ is $\eta$-isolated for any $\eta>4$ if we initialize Alg. 23.1 at a sufficiently large $y$.
This compression queries evaluation oracle twice while performing Schröder’s single iteration for computing $z$ and otherwise runs at the cost of performing Alg. 9.1, estimated in Cor. 9.1. We can decrease the latter cost bound for a large class of polynomials $p$ by choosing larger tolerance $\epsilon$. Then we will yield desired compression if the algorithm outputs $r$ exceeding $\epsilon/4$. Otherwise we can recursively repeat the computations for tolerance recursively decreased to $\epsilon/4$.
@@ 23.3 Extension to any well-isolated disc
Now suppose that #$\#(D(0,1))=m<d$ and $i(D(0,1))\geq\theta$ for a sufficiently large $\theta$ of order $d^{O(1)}$. Then verify that the $d-m$ external roots, lying outside the disc $D(0,1)$, little affect the proof of Thm. 23.1 and our whole study in the previous subsection.
Thus we first perform $O(\log(d))$ subdivision steps (see Observation 20.2). We do not need compression if these steps break the component into more than one component or compress it into an $\epsilon$-disc. Otherwise the component remains compact and hence in $O(\log(d))$ subdivision steps shrinks enough to become $d^{\nu}$-isolated for any fixed constant $\nu$, in which case we apply our compression algorithm.
23.4 A fast root-counter and a fast e/i test based on Schröder’s iterations
@@
Then our study in this subsection implies that the iteration outputs a point $z=z(m)$ lying in or near the disc $D(0,1)$, say, in the disc $D(0,2)$, if and only if #$\#(D(0,1))=m$, that is, we count roots in $D(0,1)$ at the cost of evaluation of $p(x)$ at two points.
@@
In this case $|m~{}{\rm NR}(y)-y|\leq 2$, and so $|{\rm NR}(y)-y|/m\leq 2/m$.
If #$\#(D(0,1))=\bar{m}\neq m$, then the iteration outputs a point $z=z(m)$ that lies at the distance at least $|\bar{m}-m|\cdot|{\rm NR}(y)|-2$ from the disc $D(0,2)$, which must be large if $|y|$ is sufficiently (although reasonably) large.
Extension to a fast e/i test. Now suppose that the disc $D(0,1)$ is sufficiently well isolated and apply this root-counter to the polynomial $xp(x)$ for a large $y$ and $m=1$, in which case Schröder’s iterations turn into Newton’s. In this case root-counter will show that the disc $D(0,1)$ contains no roots of $p(x)$ if and only if its output lies in that disc.
@@ 23.5 Overall complexity of our root-finders
We have reduced our root-finding in an isolated disc to performing at most $12m-$ 1 compression steps and in addition $O(m)$ e/i tests. We have also reduced a compression step to (i) performing $O(\log(d))$ additional subdivision steps (included to strengthen isolation of input discs for compression), (ii) application of Schröder’s single iteration, which evaluates polynomial $p(x)$ at two points, and (iii) performing $O(\log(b))$ times $m$-tests for approximation of the root radius $r_{d-m+1}(z,p)$.
Overall our root-finders apply e/i tests $O(m\log(d))$ times, $m$-tests $O(m\log(b))$ times, and at most $22m-$ 2 Schröder’s iterations. Clearly, the computational cost of Schröder’s is dominated, while in Sec. 15 we reduced an $m$-test is to $O(1)$ e/i tests, specified in Sec. 14. Based on the latter results we can strengthen our estimates of Sec. 18 and arrive at the complexity bounds of our Main Theorems. Alternatively, instead of estimating the cost of e/i tests and reduction to them of $m$- tests, we can use our estimates for the complexity of $\ell$- tests of Part II for $\ell=1$ and $\ell=m$ and would still arrive at our Main Theorems 2 – 4.
Part V: Factorization, Root Isolation, and Functional Iterations
@@
24 Polynomial Factorization, its links to Root-finding, and some bit-complexity estimates
24.1 Polynomial Factorization and its links to Root-finding (the results)
@@
The root-finders of [134, 96, 103] solve Problems 0 and 0 by extending the the solution of the following problem of independent importance:
Problem F. Approximate Polynomial Factorization: [^35] The solution of Problem F can be also extended to isolation of the zeros of a square-free polynomial with integer coefficients and approximation of matrix eigenvalues (see our Remark 25.4 and Sec. 11) and has various other applications to modern computations, e.g., to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147]. For a polynomial of degree $d$ the known sharp upper and lower bounds on the input precision and number of bit operations for root-finding exceed those for numerical factorization by a factor of $d$. Given a real $b$ and $d+1$ coefficients $p_{0},p_{1},\dots,p_{d}$ of (1.1), compute $2d$ complex numbers $u_{j},v_{j}$ for $j=1,\dots,d$ such that
$|p-\prod_{j=1}^{d}(u_{j}x-v_{j})|\leq\epsilon|p|~{}{\rm for}~{}\epsilon=1/2^{b}.$ (24.1)
@@
Definition 24.1.
Polynomials with bounded root cluster sizes. The zero set of $p$ consists of clusters, each made up of a constant number of zeros and having a diameter at most $1/2^{3-b/d}$; furthermore, these clusters are pairwise separated by the distances at least $1/d^{\tilde{O}(1)}$.
@@
In Sec. 24.3 we reduce Problem 0 to Problem F and vice versa by applying Schönhage’s results of [134, Cor. 4.3 and Sec. 19]:
Theorem 24.2.
(i) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}(bd,d))$ for any polynomial $p(x)$ of degree $d$, (ii) $\mathbb{B}_{F}(b,d)=\tilde{O}(\mathbb{B}_{0}(b+d+\log(d),d))$ for any polynomial $p(x)$ of degree $d$, and (iii) $\mathbb{B}_{0}(b,d)=\tilde{O}(\mathbb{B}_{F}2(b,d))$ for polynomials of Def. 24.1, with bounded cluster sizes.
By combining Thms. 24.1 and 24.2 obtain the following estimates.
Theorem 24.3.
@@
(ii) apply the algorithm that supports Thm. 24.3 to the $m$th degree factor of the polynomial $\prod_{j=1}^{d}(u_{j}x-v_{j})$ of (24.1) that shares with this polynomial all its zeros lying in the disc $D(1+\beta)$ for positive $\beta$ of Def. 1.1.363636We can slightly speed up stage (i) because at that stage we actually only need to approximate factorization of the factor $f$ of degree $m$ of $p$ that shares with $p$ its zeros in $D$.
24.2 Lower bounds on the bit operation complexity